. Appendix B Specimen Financial Statements: The Procter & Gamble Company Once each year, a corporation communicates to its st2016 2017 $ 65,058 32,535 18,568 $ 65,299 32,909 18,949 2015 $ 70,749 37,056 20,616 2,028 13,441 579 182 11,049 626 149 440 32016 2015 2017 S 15,411 $ 10,604 $ 7,144 Consolidated Statements of Comprehensive Income Amounts in millions; Years ended Jun2,139 Prepaid expenses and other current assets Current assets held for sale TOTAL CURRENT ASSETS PROPERTY, PLANT AND EQUIPMEConsolidated Statements of Shareholders Equity Accumulated Other Total Common Additional Reserve for Comprehensive Non Share(85) (85) $642 $57,983 85 15,411 1,275 Noncontrolling interest, net BALANCE JUNE 30, 2016 2,668,074 $4,009 $1,038 $63,714 ($110,604 3,078 7,144 3,134 15,411 2,820 543 351 (601) (5,490) 335 (815) (41) (322) 71 (149) (43) 450 35 116 1,285 204 184 337 (- Pre-uivestiture audition or restricted casri related to me Beauty branius civestilure (074) 1990) Cash transferred at closiThe financial statements of P&G are presented in Appendix B. The companys complete annual report, including the notes to theWhat were P&Gs cash flows from its operating, investing, and financing activities for 2017? What were its trends in net cash(Round answers to 1 decimal place, e.g. 52.7.) P&Gs net cash provided by operating activities by % from 2015 to 2016. P&Gs

Please show work as well if you can, thank you.
Answer #1

Working Capital = Current Assets - Current liabilities

Working capital

2017 = 26494 - 30210 = - 3716

2016 = 33782 - 30770 = 3012

Net cash from oprating

2017 = 12753

2016 = 15435

2015 = 14608

Net cash from investing

2017 = -5689

2016 = -5575

Net cash from financing

2017 = -8568

2016 = -9213

From 2015 to 2016 net cash from operating increased by:

(15435 - 14608)/14608 = 5.7%

From 2016 to 2017 net cash from operating increased by:

(12753-15435)/15435 = -17.4%

Current Cash Debt Coverage ratio

Current Cash Debt Coverage ratio =  net Cash provided by operating activities ÷ Average total current liabilities

where,

Average total current liabilities =  (current liabilities of previous year + current liabilities of this year) ÷ 2

For 2017,

Average total current liabilities = (30770 + 30210)/2 = 30490

Current Cash Debt Coverage ratio = 12753 / 30490 = 0.42

Cash Debt Coverage ratio

Cash Debt Coverage ratio = net Cash provided by operating activities ÷ total liabilities

total liabilities for 2017 = 64628

Cash Debt Coverage ratio = 12753 / 64628

Cash Debt Coverage ratio = 0.2

Free Cash FLow

Free Cash FLow =

Factor Location + Cash Flow from Operating Activities Statement of Cash Flows + Interest Expense Income Statement - Tax Shiel

Tax rate for the company = income taxes / earning from continous operation before income taxes    (Income statement)

Tax rate for the company = 3063/13257 = 23.1%

Tax shield on interest expense = Tax rate as calculated above x Net interest expense

Tax shield on interest expense = 23.1% x 404 = 93.34

Free Cash FLow = 12753 + 404 - 93.34 - 3384

Free Cash FLow = 9679.66